OpenStudy (anonymous):

Can anybody explain to me the normal (Gaussian) equation?Can anybody explain to me the normal (Gaussian) equation?@Mathematics

OpenStudy (anonymous):

\[P(x) = \frac{1}{\sqrt{2\pi \sigma ^2}}* \exp(-\frac{(x-\mu)^2}{2\sigma ^2})\]